OpenStudy (anonymous):

State the horizontal asymptote of the rational function. f(x)=(5x+1)/(9x-1)

OpenStudy (anonymous):

@satellite73 @ZakaullahUET

OpenStudy (anonymous):

can u point the denominator of ur equation

OpenStudy (anonymous):

it is 9x-1 equate it to zero what u get is ur answer 9x-1=0 x=1/9

OpenStudy (anonymous):

the only answer choices are None y = 3/2 y = 5/9 y = 0

OpenStudy (anonymous):

y = 5/9 is the answer sorry i figured out wrong above leme check that again but ur answer is 5/9

OpenStudy (anonymous):

ok thankyou